OpenStudy (anonymous):

2 questions on Polynomials. Can i get some help???

OpenStudy (anonymous):

please ask :)

OpenStudy (anonymous):

There

OpenStudy (anonymous):

first one: (x+5)(2x-4)=2x^2+6x-20

OpenStudy (anonymous):

second one: (x-3)^2+(x-3)^2/4+(x-3)^2/16+....=(x-3)^2*(1+1/4+1/16....)=4*(x-3)^2/3 ok??

OpenStudy (anonymous):

Yes :)